a total of 15000 is invested in two funds paying 4%and 5% simple interest. the annual interest is $700. how much is invested in each fund?
Answer by mananth(16949)

total investment = $15000
4% --------------- $x
5%-----------------$15000-x
..
interest = $700
...
0.04x+0.05(15000-x)=700
0.04x+750-0.05x=700
-0.01x=-50
x= -50/-0.01
x= 5000 at 4%
balance 10000 at 5%
...
m.ananth@hotmail.ca